Given that 2i is a root of the polynomial equation x^4 - 45x^2 - 196 = 0,find the solution set for the equation.
Answer by Alan3354(69443)

Given that 2i is a root of the polynomial equation x^4 - 45x^2 - 196 = 0,find the solution set for the equation.
------------
Complex roots occur in conjugate pairs, so -2i is also a root.
(x+2i)*(x-2i) = x^2 + 4
Divide the polynomial by (x^2 + 4) and solve the remaining quadratic.